# Asymptotic expansions for solutions of differential equations having a coalescing turning point and double pole, with an application to Legendre functions

### v3

@@ Abstract
The asymptotic behavior of solutions to the second-order linear differential equation $d^{2}w/dz^{2}={u^{2}f(\alpha,z)+g(z)}w$ is analyzed for a large real parameter $u$ and $\alpha\in[0,\alpha_{0}]$, where $\alpha_{0}>0$ is fixed. The independent variable $z$ ranges over a complex domain $Z$ (possibly unbounded) on which $f(\alpha,z)$ and $g(z)$ are analytic except at $z=0$, where the differential equation has a regular singular point. For $\alpha>0$, the function $f(\alpha,z)$ has a double pole at $z=0$ and a simple zero in $Z$, and as $\alpha\to 0$ the turning point coalesces with the pole. Bessel function approximations are constructed for large $u$ involving asymptotic expansions that are uniformly valid for $z\in Z$ and $\alpha\in[0,\alpha_{0}]$. The expansion coefficients are generated by simple recursions, and explicit error bounds are obtained that simplify earlier results. As an application, uniform asymptotic expansions are derived for associated Legendre functions of large degree $\nu$, valid for complex $z$ in an unbounded domain and for order $\mu\in[0,\nu(1-\delta)]$, where $\delta>0$ is arbitrary.
@@ 1 Introduction
with a regular singularity at $z=0$ and, for $0<\alpha\leq\alpha_{0}$, a simple turning point, that is, a zero of $f(\alpha,z)$ at $z=x_{t}\in(0,\infty)$. The location of this turning point is a continuous real function of $\alpha$ and tends to 0 as $\alpha\to 0$.
@@
Hence, the regular singularity at $z=0$ has exponents $\pm\frac{1}{2}\mu+\frac{1}{2}$, where
@@
In the case $\alpha=0$ we assume that $f(0,z)$ is nonzero, has a simple pole at $z=0$, and is negative (respectively positive) for positive (respectively negative) $z\in D$. We construct asymptotic expansions with explicit error bounds in a domain $Z\subseteq D$ that contains both the pole and the turning point, with validity uniform for $0\leq\alpha\leq\alpha_{0}$. For historical background on coalescing turning points and double poles, see [1].
We shall apply our theory to obtain uniform asymptotic approximations for associated Legendre functions; our general theory is also applicable to Laguerre polynomials [8]; further applications include spheroidal wave functions, Whittaker’s confluent hypergeometric functions, and Jacobi, ultraspherical, and Charlier polynomials.
Let us introduce the well-known Liouville-Green (LG) variables (see [9, Chap. 10, Eq. (2.02)])
@@
Classical expansions are derived from Eq. 6 in terms of elementary functions. We use the form in which an asymptotic series appears in the exponent of an exponential [3], rather than the more traditional form in which an expansion multiplies an exponential function of argument $\pm u\xi$ [9, Chap. 10].
@@
To obtain expansions valid at both $z=0$ and $z=z_{t}$, following [1, Eq. (2.1)], we introduce new independent and dependent variables $\zeta$ and $\tilde{W}$ via
@@
In [1] asymptotic solutions of Eq. 16 were obtained in the form
@@
    $\xi$ lies on a continuous directed path $\mathcal{L}$, consisting of a finite chain of $R_{2}$ arcs (as defined in [9, Chap. 5, §3.3]), that runs from $\xi_{1}=\infty+\alpha\pi i/2$ to $\xi_{2}=-\infty+\alpha\pi i/2$, with points $v\in\mathcal{L}$ satisfying (i)–(ii) and $\Re(v)$ decreasing along the path.
@@ 2 LG solutions
First, following [3, Eqs. (1.11) and (1.15)] we define
@@
Then, from [3, Eq. (1.14)] we have for the odd LG coefficients
@@
Equation Eq. 25 can also be used for the even LG coefficients $E_{2s}(\alpha,z)$ ($s=1,2,3,\ldots$). However, as shown in [7] using Abel’s identity, these coefficients can be determined without integration by expanding the formal identity
@@
From [3, Eqs. (1.16)-(1.18)], denote the fundamental LG solutions of Eq. 16, which for convenience we consider as functions of $z$, by
@@
These bounds hold where the integrals converge; thus, the integration paths must avoid the turning point $z=z_{t}$, since, by Eqs. 7, 23, and 24, the coefficients are not integrable at this point. At the singularities $z=z_{1,2}$ they are assumed to converge. The conditions on $f(\alpha,z)$ and $g(z)$ to ensure this are well known, for example given by [9, Chap. 10, Thm. 4.1] when they are poles, and [9, Chap. 10, Exercise 4.1] when they are at infinity.
@@ 3 LG expansions for Bessel functions
From Bessel’s equation (see, for example, [2, Eq. 10.2.1]), the scaled Bessel functions $\zeta^{1/2}\mathcal{C}_{\mu}(u\zeta^{1/2})$ ($\mathcal{C}=J,\,H^{(1)},\,H^{(2)}$) are readily verified to be solutions of
@@
The LG coefficients are derived as in Section 2 (see [3]), with $z$ replaced by $\zeta$ and $f,\,g,\,\psi$ replaced by $\hat{f},\,\hat{g},\,\hat{\psi}$. To facilitate the integrations, define
@@
Next, from the well-known behavior of the Hankel functions at infinity [2, Eqs. 10.17.5 and 10.17.6], we have
@@
Error bounds are obtained similarly to Eqs. 35, 36, and 37. Using Eq. 69 together with [3, Thm. 1.1], we have
@@
We proceed analogously to Eqs. 61 and 62, using [2, Eqs. 10.17.11 and 10.17.12]
@@ 4 Main expansions
Next, we solve the pair of equations for Eqs. 76 and 77 for $A_{\mu}(u,\alpha,z)$ and $B_{\mu}(u,\alpha,z)$, with the aid of the Hankel function Wronskian [2, Eq. 10.5.5]. This yields the explicit representations
@@
As noted in [6, Remark 1] for similar expansions, $\mathrm{A}_{s}(\alpha,z)=\tilde{q}_{2s}(\alpha,z)$ ($s=1,2,3,\ldots$) and $\mathrm{B}_{s}(\alpha,z)=\hat{\beta}^{-1}q_{2s+1}(\alpha,z)$ ($s=0,1,2,\ldots$), where for $s=1,2,3,\ldots$
@@
To do so, we first prove that $A_{\mu}(u,\alpha,z)$ and $B_{\mu}(u,\alpha,z)$ are analytic in $Z$. To this end, using Eqs. 39, 41, 76, and 77 and [2, Eqs. 10.4.4 and 10.4.6], we express the Frobenius solutions $\tilde{W}_{\pm\mu}(u,\alpha,z)$ similarly as
@@
(90) $B_{\mu}(u,\alpha,z)=\frac{\pi u}{4\zeta^{1/2}\sin(\mu\pi)}\left\{\tilde{W}_{\mu}^{(0)}(u,\alpha,z)J_{\mu}\left(u\zeta^{1/2}\right)\right.\\ \left.-\tilde{W}_{-\mu}(u,\alpha,z)J_{\mu}\left(u\zeta^{1/2}\right)\right\}.$
Expanding the Frobenius series for all four functions on the RHS of Eqs. 89 and 90 about $z=\zeta=0$ (see Eq. 38 and [2, Eq. 10.2.]) shows that $A_{\mu}(u,\alpha,z)$ and $B_{\mu}(u,\alpha,z)$ are analytic at this point for noninteger $\mu$; for integer $\mu$, this also holds by taking limits of the resulting Maclaurin series. Analyticity elsewhere for $z\in Z$ follows from analytic continuation.
@@
We now use the following theorem. Its proof is a routine modification of that of [4, Thm. 3.1] for a single isolated singularity; see also [5, Thm. 4.2].
@@ 5 Associated Legendre functions
Standard solutions are given by $y(u,\alpha,t)=L_{\nu}^{\pm\mu}(t)$ where $L=\mathrm{P},\mathrm{Q},P,Q$ (see [2, §14.3]). For simplicity, we restrict our attention to the half-plane $|\arg(t)|\leq\pi/2$, with extensions to other parts of the complex plane achievable through appropriate connection formulas; see, for example, [2, §14.24].
@@
Next, following [5, Eq. (5.11)], define
@@
From this we note that [2, Eq. 14.9.14] $\mathsf{Q}^{-\mu}_{\nu,\pm 1}(t)=\mathsf{Q}^{\mu}_{\nu,\pm 1}(t)$. The significance of these solutions is that they are analytic in the $z$ plane having cuts along $(-\infty,-1]$ and $1,\infty)$, with $\mathsf{Q}^{\mu}_{\nu,\pm 1}(t)$ recessive at $t=\infty\pm i0$ ($z=-\infty\mp i0$).
From [2, Eqs. 14.3.1 and 14.8.1] the solution of Eq. 97 that is recessive at $t=1$ is the Ferrers function given by
@@
where $\mathbf{F}$ is the scaled hypergeometric function defined by [2, Eq. 15.2.2]. In this, we can replace $\mu\neq 1,2,3,\ldots$ by $\mu$ to obtain the solution of Eq. 97 that has the exponent $\mu/2$ at the singularity $t=1$. From Eq. 122 [2, Eq. 14.23.2] the following connection formula will be utilized
@@
We now match solutions of Eq. 97 that are recessive at $t=\infty\pm i0$ with the LG expansions given by Eqs. 30 and 31. To this end, from Eq. 108, Eq. 122 and [2, Eq. 14.8.15] for nonnegative $\nu$
@@
(129) $\mathsf{P}^{-\mu}_{\nu}(t)=\sqrt{\frac{2}{\pi}}\left(\frac{\sigma}{1+\alpha}\right)^{\mu}\frac{\Gamma\left(u-\mu+\tfrac{1}{2}\right)}{\sigma^{u}\Gamma(u+1)}\left(\frac{\zeta-\alpha^{2}}{\sigma^{2}-t^{2}}\right)^{1/4}\\ \times\left\{J_{\mu}\left(u\zeta^{1/2}\right)A_{\mu}(u,\alpha,z)+\frac{\zeta}{u}\frac{\partial J_{\mu}\left(u\zeta^{1/2}\right)}{\partial\zeta}B_{\mu}(u,\alpha,z)\right\},$
@@
From [5, Eq. (5.14)] the numerically satisfactory companion Ferrers function to $\mathsf{P}_{\nu}^{-\mu}(t)$ for $0\leq t<1$ is the real-valued function given by
@@
Consequently, on referring to [2, Eq. 10.4.4], we find the following representation for $0\leq t<1$
@@
Finally, consider solutions to Eq. 97 that are real for $t=x\in(1,\infty)$ (thus, from Eqs. 12 and 111 z=1xz=1-x and $\zeta$, both lie in $(-\infty,0)$). For the one that is recessive at $x=1$ we have from [2, Eq. 14.23.4]
@@
Now $\zeta^{1/2}=-i|\zeta|^{1/2}$ for $t=x+i0$ and $1<x<\infty$. Thus, from Eq. 129 and [2, Eq. 10.27.6]
@@
Using Eq. 122, Eq. 128 and [2, Eq. 10.27.8], for the solution that is real-valued for $x\in(1,\infty)$ and recessive at $x=\infty$, we similarly obtain

### v4

@@ Abstract
The asymptotic behavior of solutions to the second-order linear differential equation $d^{2}w/dz^{2}={u^{2}f(\alpha,z)+g(z)}w$ is analyzed for a large real parameter $u$ and $\alpha\in[0,\alpha_{0}]$, where $\alpha_{0}>0$ is fixed. The independent variable $z$ ranges over a complex domain $Z$ (possibly unbounded) on which $f(\alpha,z)$ and $g(z)$ are analytic except at $z=0$, where the differential equation has a regular singular point. For $\alpha>0$, the function $f(\alpha,z)$ has a double pole at $z=0$ and a simple zero in $Z$, and as $\alpha\to 0$ the turning point coalesces with the pole. Bessel-function approximations are constructed for large $u$ involving asymptotic expansions that are uniformly valid for $z\in Z$ and $\alpha\in[0,\alpha_{0}]$. The expansion coefficients are generated by simple recursions, and explicit error bounds are obtained that simplify earlier results. As an application, uniform asymptotic expansions are derived for associated Legendre functions of large degree $\nu$, valid for complex $z$ in an unbounded domain and for order $\mu\in[0,\nu(1-\delta)]$, where $\delta>0$ is arbitrary.
@@ 1 Introduction
with a regular singularity at $z=0$ and, for $0<\alpha\leq\alpha_{0}$, a simple turning point, that is, a zero of $f(\alpha,z)$ at $z=z_{t}\in(0,\infty)$. The location of this turning point is a continuous real function of $\alpha$ and tends to 0 as $\alpha\to 0$.
@@
Hence, the regular singularity at $z=0$ has exponents $\frac{1}{2}\pm\frac{1}{2}\mu$, where
@@
In the case $\alpha=0$ we assume that $f(0,z)$ is nonzero, has a simple pole at $z=0$, and is negative (respectively positive) for positive (respectively negative) $z\in D$. We construct asymptotic expansions with explicit error bounds in a domain $Z\subseteq D$ that contains both the pole and the turning point, with validity uniform for $0\leq\alpha\leq\alpha_{0}$. For historical background on coalescing turning points and double poles, see [Boyd:1986:TPS].
We shall apply our theory to obtain uniform asymptotic approximations for associated Legendre functions. There have been numerous numerical and asymptotic studies of associated Legendre functions for large parameters, both through diff eq and integral and difference approaches; see for example [Bakaleinikov:2020:UAL, Frenzen:1990:EBU, Gil:2000:CTF, Nemes:2020:LPA, Olver:1975:LFW, Olver:1997:ASF, Shivakumar:1988:EBU, Temme:2015:AMF, Tumarkin:1959:ASO, Ursell:1984:ILP, Wang:2012:ASY]. In contrast, the present results, together with the complementary expansions in [Dunster:2026:SUA], are valid when one or both parameters are large, with coefficients that are readily computable.
Our general asymptotic theory is also applicable to Laguerre polynomials, spheroidal wave functions, Whittaker’s confluent hypergeometric functions, and Jacobi, ultraspherical, and Charlier polynomials. These special functions arise in a wide range of applications, including the scattering of particles with spin where Legendre and Jacobi (rotation) functions are employed [Durand:2019:ABE], the description of dispersive shock waves in the focusing nonlinear Schrödinger equation via generalized Laguerre polynomials [Kotlyarov:2019:DSW], constructions of polynomial families (including Laguerre-type and higher-order Hermite polynomials) arising in Burgers-like nonlinear PDEs [Dattoli:2024:HHO], signal-processing frame expansions of band-limited signals using prolate spheroidal wave functions [Hogan:2016:STS], bound-state solution methods for the Schrödinger equation via Laplace transforms of the confluent hypergeometric equation [Nogueira:2016:LCB], and numerical schemes for nonlinear stochastic differential equations (driven by fractional Brownian motion) based on shifted Jacobi polynomials [Singh:NSB:2023].
We begin by introducing the well-known Liouville-Green (LG) variables (see [Olver:1997:ASF, Chap. 10, Eq. (2.02)])
@@
Classical expansions are derived from Eq. 6 in terms of elementary functions. We use the form in which an asymptotic series appears in the exponent of an exponential [Dunster:2020:LGE], rather than the more traditional form in which an expansion multiplies an exponential function of argument $\pm u\xi$ [Olver:1997:ASF, Chap. 10].
@@
To obtain expansions valid at both $z=0$ and $z=z_{t}$, following [Boyd:1986:TPS, Eq. (2.1)], we introduce new independent and dependent variables $\zeta$ and $\tilde{W}$ via
@@
In [Boyd:1986:TPS] asymptotic solutions of Eq. 16 were obtained in the form
@@ 2 LG solutions
    $\xi$ lies on a continuous directed path $\mathcal{L}$, consisting of a finite chain of $R_{2}$ arcs (as defined in [Olver:1997:ASF, Chap. 5, §3.3]), that runs from $\xi_{1}=\infty+\alpha\pi i/2$ to $\xi_{2}=-\infty+\alpha\pi i/2$, with points $v\in\mathcal{L}$ satisfying (i)–(ii) and $\Re(v)$ decreasing along the path.
@@
First, following [Dunster:2020:LGE, Eqs. (1.11) and (1.15)] we define
@@
Then, from [Dunster:2020:LGE, Eq. (1.14)] we have for the odd LG coefficients
@@
Equation Eq. 25 can also be used for the even LG coefficients $E_{2s}(\alpha,z)$ ($s=1,2,3,\ldots$). However, as shown in [Dunster:2017:COA] using Abel’s identity, these coefficients can be determined without integration by expanding the formal identity
@@
From [Dunster:2020:LGE, Eqs. (1.16)-(1.18)], denote the fundamental LG solutions of Eq. 16, which for convenience we consider as functions of $z$, by
@@
These bounds hold where the integrals converge; thus, the integration paths must avoid the turning point $z=z_{t}$, since, by Eqs. 7, 23, and 24, the coefficients are not integrable at this point. At the singularities $z=z_{1,2}$ they are assumed to converge. The conditions on $f(\alpha,z)$ and $g(z)$ to ensure this are well known, for example given by [Olver:1997:ASF, Chap. 10, Thm. 4.1] when they are poles, and [Olver:1997:ASF, Chap. 10, Exercise 4.1] when they are at infinity.
@@ 3 LG expansions for Bessel functions
From Bessel’s equation (see, for example, [NIST:DLMF, Eq. 10.2.1]), the scaled Bessel functions $\zeta^{1/2}\mathcal{C}_{\mu}(u\zeta^{1/2})$ ($\mathcal{C}=J,\,H^{(1)},\,H^{(2)}$) are readily verified to be solutions of
@@
The LG coefficients are derived as in Section 2 (see [Dunster:2020:LGE]), with $z$ replaced by $\zeta$ and $f,\,g,\,\psi$ replaced by $\hat{f},\,\hat{g},\,\hat{\psi}$. To facilitate the integrations, define
@@
Next, from the well-known behavior of the Hankel functions at infinity [NIST:DLMF, Eqs. 10.17.5 and 10.17.6], we have
@@
Error bounds are obtained similarly to Eqs. 35, 36, and 37. Using Eq. 53 together with [Dunster:2020:LGE, Thm. 1.1], we have
@@
We proceed analogously to Eqs. 61 and 62, using [NIST:DLMF, Eqs. 10.17.11 and 10.17.12]
@@ 4 Main expansions
Next, we solve the pair of equations for Eqs. 76 and 77 for $A_{\mu}(u,\alpha,z)$ and $B_{\mu}(u,\alpha,z)$, with the aid of the Hankel function Wronskian [NIST:DLMF, Eq. 10.5.5]. This yields the explicit representations
@@
As noted in [Dunster:2025:SAR, Remark 1] for similar expansions, $\mathrm{A}_{s}(\alpha,z)=\tilde{q}_{2s}(\alpha,z)$ ($s=1,2,3,\ldots$) and $\mathrm{B}_{s}(\alpha,z)=\hat{\beta}^{-1}q_{2s+1}(\alpha,z)$ ($s=0,1,2,\ldots$), where for $s=1,2,3,\ldots$
@@
To do so, we first prove that $A_{\mu}(u,\alpha,z)$ and $B_{\mu}(u,\alpha,z)$ are analytic in $Z$. To this end, using Eqs. 39, 41, 76, and 77 and [NIST:DLMF, Eqs. 10.4.4 and 10.4.6], we express the Frobenius solutions $\tilde{W}_{\pm\mu}(u,\alpha,z)$ similarly as
@@
(90) $B_{\mu}(u,\alpha,z)=\frac{\pi u}{4\zeta^{1/2}\sin(\mu\pi)}\left\{\tilde{W}_{\mu}^{(0)}(u,\alpha,z)J_{-\mu}\left(u\zeta^{1/2}\right)\right.\\ \left.-\tilde{W}_{-\mu}(u,\alpha,z)J_{\mu}\left(u\zeta^{1/2}\right)\right\}.$
Expanding the Frobenius series for all four functions on the RHS of Eqs. 89 and 90 about $z=\zeta=0$ (see Eq. 38 and [NIST:DLMF, Eq. 10.2.]) shows that $A_{\mu}(u,\alpha,z)$ and $B_{\mu}(u,\alpha,z)$ are analytic at this point for noninteger $\mu$; for integer $\mu$, this also holds by taking limits of the resulting Maclaurin series. Analyticity elsewhere for $z\in Z$ follows from analytic continuation.
@@
We now use the following theorem. Its proof is a routine modification of that of [Dunster:2021:NKF, Thm. 3.1] for a single isolated singularity; see also [Dunster:2025:LCT, Thm. 4.2].
@@ 5 Associated Legendre functions
Standard solutions are given by $y(u,\alpha,t)=L_{\nu}^{\pm\mu}(t)$ where $L=\mathrm{P},\mathrm{Q},P,Q$ (see [NIST:DLMF, §14.3]). For simplicity, we restrict our attention to the half-plane $|\arg(t)|\leq\pi/2$, with extensions to other parts of the complex plane achievable through appropriate connection formulas; see, for example, [NIST:DLMF, §14.24].
@@
Next, following [Dunster:2025:LCT, Eq. (5.11)], define
@@
From this we note that [NIST:DLMF, Eq. 14.9.14] $\mathsf{Q}^{-\mu}_{\nu,\pm 1}(t)=\mathsf{Q}^{\mu}_{\nu,\pm 1}(t)$. The significance of these solutions is that they are analytic in the $z$ plane having cuts along $(-\infty,-1]$ and $1,\infty)$, with $\mathsf{Q}^{\mu}_{\nu,\pm 1}(t)$ recessive at $t=\infty\pm i0$ ($z=-\infty\mp i0$).
From [NIST:DLMF, Eqs. 14.3.1 and 14.8.1] the solution of Eq. 97 that is recessive at $t=1$ is the Ferrers function given by
@@
where $\mathbf{F}$ is the scaled hypergeometric function defined by [NIST:DLMF, Eq. 15.2.2]. In this, we can replace $\mu\neq 1,2,3,\ldots$ by $\mu$ to obtain the solution of Eq. 97 that has the exponent $\mu/2$ at the singularity $t=1$. From Eq. 122 [NIST:DLMF, Eq. 14.23.2] the following connection formula will be utilized
@@
We now match solutions of Eq. 97 that are recessive at $t=\infty\pm i0$ with the LG expansions given by Eqs. 30 and 31. To this end, from Eq. 108, Eq. 122 and [NIST:DLMF, Eq. 14.8.15] for nonnegative $\nu$
@@
(129) $\mathsf{P}^{-\mu}_{\nu}(t)=\sqrt{\frac{2}{\pi}}\left(\frac{\sigma}{1+\alpha}\right)^{\mu}\frac{\Gamma\left(u-\mu+\tfrac{1}{2}\right)}{\sigma^{u}\Gamma(u+1)}\left(\frac{\zeta-\alpha^{2}}{\sigma^{2}-t^{2}}\right)^{1/4}\\ \times\left\{J_{\mu}\left(u\zeta^{1/2}\right)A_{\mu}(u,\alpha,z)+\frac{\zeta}{u}\frac{\partial J_{\mu}\left(u\zeta^{1/2}\right)}{\partial\zeta}B_{\mu}(u,\alpha,z)\right\}.$
@@
From [Dunster:2025:LCT, Eq. (5.14)] the numerically satisfactory companion Ferrers function to $\mathsf{P}_{\nu}^{-\mu}(t)$ for $0\leq t<1$ is the real-valued function given by
@@
Consequently, on referring to [NIST:DLMF, Eq. 10.4.4], we find the following representation for $0\leq t<1$
@@
Finally, consider solutions to Eq. 97 that are real for $t=x\in(1,\infty)$ (thus, from Eqs. 12 and 111 z=1xz=1-x and $\zeta$, both lie in $(-\infty,0)$). For the one that is recessive at $x=1$ we have from [NIST:DLMF, Eq. 14.23.4]
@@
Now $\zeta^{1/2}=-i|\zeta|^{1/2}$ for $t=x+i0$ and $1<x<\infty$. Thus, from Eq. 129 and [NIST:DLMF, Eq. 10.27.6]
@@
Using Eq. 122, Eq. 128 and [NIST:DLMF, Eq. 10.27.8], for the solution that is real-valued for $x\in(1,\infty)$ and recessive at $x=\infty$, we similarly obtain